In a particular species of plant, tall is dominant to short, and orange petals are dominant to the recessive white colour. Use T and t to symbolize the alleles for height, and F and f to symbolize the alleles for flower colour. A homozygous tall white flower is crossed with a flower heterozygous for both traits.

1.1 List the genotypes of the parents.

1.2 What are the F1 GENOTYPIC and PHENOTYPIC ratios?

1.3 What are the F2 genotypic and phenotypic ratios?
1
Expert's answer
2021-09-28T20:24:01-0400

1.1 The parents genotypes are

TTff x TtFf


1.2 The F1 genotype will be

TTFf, TTff, TtFf, Ttff

The phenotypic ratio of tall and orange colored is "\frac{4}{8}=\frac{1}{2}"

The phenotypic ratio for tall and white is "\frac{4}{8}=\frac{1}{2}"

The genotypic ratio for each is "\frac{1}{4}"


1.3 The F2 genotypic and phenotypic will be

The phenotypic ratio of tall and orange colored is "\frac{4}{8}=\frac{1}{2}"

The phenotypic ratio for tall and white is "\frac{4}{8}=\frac{1}{2}"

The genotypic ratio for each is "\frac{3}{4}"